The diameter of a circle is 1.25 inches
What is the cross sectional area of this opening in square inches and square centimeters
1
Expert's answer
2019-01-24T10:10:07-0500

The cross sectional area of a circle

A=πr^2=(πd^2)/4

=〖3.14×(1.25 in)〗^2/4=1.23 in^2

Since 1 in=2.54 cm, we obtain

A=1.23×(2.54 cm)^2=7.91 cm^2